CHARACTERIZATION OF A WATER-HEXAFLUOROBENZENE COMPLEX USING MATRIX ISOLATION INFRARED SPECTROSCOPY

\underline{JAY C. AMICANGELO}, DANIEL G. IRWIN, CYNTHIA J. LEE, and NATALIE C. ROMANO, NANCY L. SAXTON, School of Science, Penn State Erie, Erie, PA 16563.

Matrix isolation infrared spectroscopy was used to characterize a 1:1 complex of water (H_{2}O) with hexafluorobenzene (C_{6}F_{6}). Co-deposition experiments with H_{2}O and C_{6}F_{6} were performed at 17 K using nitrogen and argon as the matrix gases. New infrared bands attributable to the H_{2}O-C_{6}F_{6} complex were observed near the asymmetric stretching, symmetric stretching, and bending vibrations of the water monomer and near the C-F and C-C stretching vibrations of the C_{6}F_{6} monomer. Identification of the new infrared bands to those of the complex were established by comparing the co-deposition spectra with the spectra of the individual monomers, by performing experiments with isotopically labeled water (D_{2}O and HDO), and by matrix annealing experiments. Theoretical calculations were also performed for the H_{2}O-C_{6}F_{6} complex using ab initio and density functional theory methods. In general, the calculations predict the H_{2}O and C_{6}F_{6} vibrational frequencies in the H_{2}O-C_{6}F_{6} complex to be shifted with respect to the H_{2}O and C_{6}F_{6} monomers by similar magnitudes as to what we observe experimentally, lending support to our assignments.